OpenStudy (anonymous):

F(t)=25,000(1.0525)^t how long does it take for the account to triple?

OpenStudy (istim):

Just sub in F(t)=3(25 000)

OpenStudy (istim):

Into the equation, and solve for t. It's rather a nice one, this question.

OpenStudy (anonymous):

Where do i sub it in at?

OpenStudy (campbell_st):

the initial amount is t= 0 which gives 25000 then 75000 = 25000(1.0525)^t 3 = (1.0525)^t now use guess and check for t... easiest method

OpenStudy (istim):

Simply F(T)

OpenStudy (istim):

Yeah, Campbell's got it.

OpenStudy (campbell_st):

t will be between 18 and 25... I'll let you find it

OpenStudy (anonymous):

I guess I am not understanding the guess and check method. It's been so long since i've had to do this stuff! lol

OpenStudy (anonymous):

Do i need to log to get rid of the exp?

OpenStudy (anonymous):

Awesome now i remember, that was it thanks for getting me on the right track!

OpenStudy (campbell_st):

pick up a calculator... enter 1.0525 to the power of 20.... is this close to 3... keep changing the power

OpenStudy (anonymous):

or just Log(3)/Log(1.0525) and gives me the exact answer. I have the answer just couldnt remember how to do the problem lol

OpenStudy (istim):

CHEATER.

OpenStudy (anonymous):

Using logarithms isnt cheating. Its actually how we were taught, im just a moron and forgot.

OpenStudy (istim):

Lol ok.

OpenStudy (istim):

Well, I guess we're good here then.

OpenStudy (anonymous):

Yup you guys got me thinking the right way so that counts.